OpenStudy (lisa123):

Find the dervivative of f (x)=-3/x at x=-4

OpenStudy (irishboy123):

\[f (x)=-\frac{3}{x} = -3 x^{-1}\] use \[\frac{d }{dx}(x^n) = n x ^{n-1} ; \ \frac{d }{dx}(ax^n) = n ax ^{n-1}\]